Ching leaves San Francisco and travels toward Los Angeles at 50 miles a hour. An hour later Phil leaves Los Angeles and travels toward San Francisco at 60 miles a hour. If the two cities are 380 miles apart, how many hours will it take for Ching to meet Phil?
Please explain!

Answer by checkley71(8403)

50T+60(T-1)=380
50T+60T-60=380
110T=380+60
T=440/110
T=4 HOURS AFTER CHING LEAVES THEY WILL MEET.
PROOF:
50*4+60*3=380
200+180=380
380=380